write the equationof a line that goes through points -4, 3 and 3, 7 in oint-slope form
Answer by mananth(16949)

x1 y1 x2 y2
-4 3 3 7

slope m = (y2-y1)/(x2-x1)
( 7 - 3 )/( 3 - -4 )
( 4 / 7 )
m= 4/7
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point(-4, 3) in
Y = m x + b
3.00 = 4/7 * -4 + b
3.00 = -16/7 + b
b= 37/7

So the equation will be
Y = 4/7 x + 37/7
m.ananth@hotmail.ca